OpenStudy (anonymous):

square root 12, 3.48, - square root 13, 10/3 order from least to greatest.

OpenStudy (compassionate):

\[\sqrt{12}, \space 3.48, \space -\sqrt{13}, \space \frac{ 10 }{ 3}\] So just find the root of 12 and -13, and 10/3, then rearange them. Tip: use a calculator.